OpenStudy (anonymous):

Find the linear approximation of the function z=ln(x-3y) at the point (7, 2). Use the linear approximation to approximate f (6.9, 2.06). Compare the approximation to the actual value at that point. What is the relative error of the approximation there? (To find the relative error a decimal calculate the difference between the actual value and the approximation and then divide by the actual value. This can be converted into a percent.) Repeat this for f (6.8, 2.07). What do you observe?

myininaya (myininaya):

\[f(x,y) \approx f(a,b)+f_x(a,b)(x-a)+f_y(a,b)(y-b) \text{ <--- linear approximation }\]

myininaya (myininaya):

Do you know how to use that formula?

OpenStudy (anonymous):

no

myininaya (myininaya):

Do you know how to take the partial of f with respect to x?

OpenStudy (anonymous):

no

myininaya (myininaya):

This is an example of how to take partials Ok what if I give you \[f(x,y)=x^2y\] \[f_x(x,y)=2xy\] \[f_y(x,y)=x^2\] So in f_x you see I treated everything like a constant except x in f(x,y) f_y , I treated everything like a constant except for y in f(x,y)

myininaya (myininaya):

You try taking the partial of our f in your problem with respect to x Remember you will have to use chain rule

myininaya (myininaya):

Can I see what you did?

myininaya (myininaya):

You should know how to take derivative of this: \[\ln(x-3)\] so how do you find derivative of this?

myininaya (myininaya):

I'm asking for you to evaluate: \[\frac{d}{dx}(\ln(x-3))\]

OpenStudy (anonymous):

1/x-3

myininaya (myininaya):

ok so if you have \[f(x,y)=\ln(x-3y) => f_x(x,y)=\frac{1}{x-3y}\]

myininaya (myininaya):

You treat y like it was a constant the derivative of inside is 1 when y is a constant right?

myininaya (myininaya):

Now you try f_y

OpenStudy (anonymous):

-3/x-3y

myininaya (myininaya):

that's right

myininaya (myininaya):

So we have \[f(x,y) \approx f(a,b)+f_x(a,b)(x-a)+f_y(a,b)(y-b) \] So (a,b)=(7,2) wherever you see a put 7 and wherever you see b put 2

myininaya (myininaya):

What is f(7,2) what is f_x(7,2) what is f_y(7,2)

OpenStudy (anonymous):

What is f(7,2)=0 what is f_x(7,2)=1 what is f_y(7,2)-=3

myininaya (myininaya):

great and the last one I think you mean =-3

OpenStudy (anonymous):

yes

myininaya (myininaya):

\[f(x,y)\approx 0+1(x-7)-3(y-2)\] and i replaced the a with 7 and the b with 2 now find f(6.9,2.06) using our approximation for f(x,y) just replace x with 6.9 and y with 2.06 and simplify

OpenStudy (anonymous):

=x-3y-1 =(6.9)-3(2.06)-1 =-0.28

myininaya (myininaya):

ok yeah the approximation i got was also -.28

myininaya (myininaya):

let me know if you need anymore help with the rest of your problem